# Cohomology of Lie algebroids over Algebraic spaces

### v6

@@ Abstract.
We consider a Lie algebroid over an algebraic space (or topological ringed space) as a quasicoherent sheaf of Lie-Rinehart algebras. We express hypercohomology for a locally free Lie algebroid (not necessarily of finite rank) as a derived functor, and simplify it via Čech cohomology. Furthermore, we define the Hochschild hypercohomology of a sheaf of generalized bialgebras and study the cases of the universal enveloping algebroid and the jet algebroid of a Lie algebroid. In the sequel, we present a version of Hochschild-Kostant-Rosenberg theorem for a locally free Lie algebroid, as well as its dual version.
@@
The author acknowledges support from the C. S. I. R. fellowship grant $2015$ during his PhD at Indian Institute of Technology Kanpur.
 [^1] AMS Mathematics Subject Classification : $16$S30303030, $16$T10,1010,10 ,$17$B56565656.
@@ 1. Introduction
The notion of Lie algebroids play a prominent role in geometry as generalized infinitesimal symmetries of spaces related to the corresponding global symmetries of spaces, called Lie groupoids [Mac05]. Lie algebroids over a $C^{\infty}$-manifold is joint generalization of tangent vector bundle over the manifold and Lie algebras [MM10]. An algebraic analogue for Lie algebroids is known as Lie-Rinehart algebras [Hue90, Kap07, Mac14, Bru17]. In the context of complex geometry and algebraic geometry, Lie algebroids over analytic spaces [Pym13] and over algebraic varieties [Kap07, CVdB10] respectively has been studied in sheaf theoretic language, which are joint generalization of the tangent sheaf and sheaves of Lie algebras [Kap07, Ram05, Pym13]. We study Lie algebroids over an algebraic space (in short we call it as $a$-spaces) or a topological ringed space $(X,\mathcal{O}_{X})$ as certain quasicoherent sheaves of Lie-Rinehart algebras [JV21, MS22], where $\mathcal{O}_{X}$ is a sub-sheaf of algebras of the sheaf of continuous functions $C_{X}^{0}$ on $X$ [Ram05, Muk15]. This combines the notion of Lie algebroids over the three types of base spaces as special cases. To study calculus on these (smooth or singular) geometric objects in a unified manner, we need the notion of Lie algebroids in algebro-geometric settings. It allows one to treat several geometric structures, such as Poisson analytic spaces, singular foliations or generalized involutive distributions (see [Fer02, Pym13]). In particular, the sheaf of logarithmic derivations $\mathcal{T}_{X}(-logY):=\{D\in\mathcal{T}_{X}~{}|~{}D(\mathcal{I})\subset\mathcal{I}\}$ for some (principal or free) divisor $(Y:=V(\mathcal{I}),~{}\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}})$ of a (smooth or complex) manifold $(X,\mathcal{O}_{X})$ with the tangent sheaf $\mathcal{T}_{X}=\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$, is one of the key object of study in this general context (see [CJNMM96, Pym13, MS22]).
@@
Associated with a Lie algebroid over an $a$-space we have two canonical generalized bialgebras, one is universal enveloping algebroid and other one is jet algebroid [MS22]. Both of them play an important role in order to study homological algebra for a Lie algebroid. The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ of a Lie algebroid $\mathcal{L}$ (generalization of sheaf of differential operators on a manifold) is sheafification of the presheaf of universal enveloping algebras of Lie-Rinehart algebras associated with each space of sections of $\mathcal{L}$ (see [Kap07, CVdB10, CRvdB10, Pym13, Bru17]). It has a canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structure [MS22], similar structures are present in [CVdB10, Kap07] by different names. Moreover the dual of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, $\mathscr{J}(\mathcal{O}_{X},\mathcal{L}):=\mathscr{H}om_{\mathcal{O}_{X}}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\mathcal{O}_{X})$ is the jet algebroid of $\mathcal{L}$ (see[CVdB10, CRvdB10, Pym13, BP15]). It is sheafification of the presheaf of jet algebras associated with the sheaf of Lie-Rinehart algebras. It has a canonical commutative, associative, unital $\mathcal{O}_{X}$-algebra structure induced from co-commutative co-associative counital coalgebra structure of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ over $\mathcal{O}_{X}$ and also have a $\mathbb{K}_{X}$-counital coalgebra structure with the canonical $\mathcal{L}$-module structure. Thus, together all the information it forms an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra [MS22].
@@
In Section 2, we recall some of the preliminary notions require for the article. In Section 3, we consider Lie algebroid hypercohomology over some special $a$-spaces. First, we recall that the Lie algebroid cohomology of a Lie algebroid $\mathcal{L}$ (locally free $\mathcal{O}_{X}$-module of finite rank) over a Noetherian separated scheme [Bru17] or over a complex manifold $(X,\mathcal{O}_{X})$ [Pym13] is expressed by derived functor $Ext$ as $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$ for some coherent $\mathcal{O}_{X}$-module $\mathcal{E}$ together with a $\mathcal{L}$-module structure. Here, we show that we can define Lie algebroid (locally free $\mathcal{O}_{X}$-module but not necessarily of finite rank) cohomology combining all of the three geometric set up (including the singular cases) and we produce an analogous result by using the result of Lie-Rinehart cohomology as derived functor for a projective Lie-Rinehart algebra [Rin63] as a local description. We apply this result for the sheaf of logarithmic derivations and for $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras. Using the cohomology groups we consider (first) Chern classes [Maa07, Maa15] in this general set up. Furthermore, we express the Lie algebroid cohomology in terms of Čech cohomology by considering a good open cover [Ram05, Ste15, BMRT15]. In Section 4, we consider algebraic (analytic) de Rham cohomologies for some free divisors associated with principal ideal sheaves [MS22] and compute the corresponding logarithmic de Rham cohomologies [CJNMM96, CM99]. We view these cohomologies as Lie algebroid (hyper)cohomology as described in [Pym13, BMRT15, Bru17].
@@
In the first part of Section 5, we define Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra. As special cases of it, we study Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ associated with a Lie algebroid $\mathcal{L}$. After that, we present a version of Hochschild-Kostant-Rosenbergh (HKR) theorem for locally free Lie algebroids (locally free $\mathcal{O}_{X}$-modules but not necessarily of finite rank) over any of the special $a$-spaces. It provides an isomorphism of graded vector spaces between Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and hypercohomology of the sheaf of $\mathcal{L}$-poly vector fields, that is $\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$. This result we get by locally using the algebraic counterpart for projective Lie-Rinehart algebras described in [KP11]. We derive a result in sheaf cohomology context to simplify the hypercohomology $\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$. Moreover, we present the HKR theorem in some of the special cases. Next we discuss about the dual version of HKR theorem in the generalize setup. Thus we show that for a locally free Lie algebroids $\mathcal{L}$ of finite rank over any of the special $a$-space $(X,\mathcal{O}_{X})$, we get a canonical isomorphism of vector spaces between Hochschild hypercohomology of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ and the Lie algebroid (hyper)cohomology of $\mathcal{L}$, that is $\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{O}_{X})$. This result we get by locally using the algebraic counterpart as described in [KP11]. Both of these hypercohomologies are obtained by the derived functor $Cotor$. At last, we apply the dual HKR theorem for the tangent sheaf $\mathcal{L}:=\mathcal{T}_{X}$ over non-singular $a$-spaces and obtain some interesting results using the (smooth, analytic or algebraic) de Rham cohomology of $X$. Note that, in [CVdB10, CRV12], to study precalculus upto homotopy for Lie algebroids of finite rank, the HKR (and its dual version) theorem appears in a more general approach. This includes an isomorphism (not exactly given by the HKR-morphism but by the HKR-morphism twisted by the square roots of the Todd genus) in the category of Gerstenhaber algebras.
@@ 2.1. Lie algebroids over algebraic spaces
Here we consider some special locally ringed spaces [Wei13], and then define Lie algebroids over that spaces [MS22].
@@
Let $X$ be a topological space and $\mathcal{O}_{X}$ a $\mathbb{K}_{X}$-subalgebra of the sheaf of continuous functions $C_{X}^{0}$ on $X$. The pair $(X,\mathcal{O}_{X})$ is said to be an algebraic space (or topological ringed space), or simply, $(X,\mathcal{O}_{X})$ is an $a$-space.
@@
Consider smooth manifolds, complex manifolds, analytic spaces and algebraic varieties, all with their associated structure sheaf, as special $a$-spaces.
@@
A Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ is a quasi-coherent sheaf of $(\mathbb{K}_{X},\mathcal{O}_{X})$-Lie-Rinehart algebras.
That is, $\mathcal{L}$ is a $\mathbb{K}_{X}$-Lie algebra and a quasi-coherent $\mathcal{O}_{X}$-module equipped with a homomorphism $\mathfrak{a}:(\mathcal{L},[\cdot,\cdot])\rightarrow(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c})$ of $\mathcal{O}_{X}$-modules and $\mathbb{K}_{X}$-Lie algebras, called the anchor map. The map $\mathfrak{a}$ satisfies the Leibniz rule:
$[D,fD^{\prime}]=f[D,D^{\prime}]+\mathfrak{a}(D)(f)D^{\prime}$ for every sections $f\in\mathcal{O}_{X}$ and $D,D^{\prime}\in\mathcal{L}$
@@
Lie algebroids over a smooth manifold $X$ is equivalent to locally free Lie algebroids of finite rank over the $a$-space $(X,C_{X}^{\infty})$.
@@
The tangent sheaf $\mathcal{T}_{X}:=\mathcal{D}er_{\mathbb{C}_{X}}(\mathcal{O}_{X})$ of an $a$-space $(X,\mathcal{O}_{X})$ has a standard Lie algebroid structure, described by the pair $(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c},id)$. In particular, when $X$ is a real smooth manifold (complex manifold), the sheaf of smooth (holomorphic) vector fields $\mathfrak{X}_{X}$ on $X$ is isomorphic to $\mathcal{T}_{X}$, forming a locally free Lie algebroid of rank equals to dimension of the manifold $X$.
The sheaf of differential $1$-forms $\Omega^{1}_{X}$ over a Poisson manifold $X$ has a canonical Lie algebroid structure. For smooth and analytic cases the details can be found in [Fer02] and [Pym13] respectively.
@@
A singular foliation $\mathcal{F}$ on a real smooth manifold or a complex manifold $(X,\mathcal{O}_{X})$ is an $\mathcal{O}_{X}$-submodule of the Lie algebroid $\mathcal{T}_{X}$ (or $\mathfrak{X}_{X})$, which is $(i)$ stable under the Lie bracket and $(ii)$ locally finitely generated. It provides a Stefan-Sussmann distribution on $X$, forms a Lie algebroid over $(X,\mathcal{O}_{X})$. In particular, regular foliations arise from involutive sub bundles of the tangent bundle (providing Frobenius distributions). See [Fer02, Pym13] for details.
@@
Let $X$ be a complex manifold and $\mathcal{O}_{X}$ be the sheaf of holomorphic functions. The vanishing set (or the zero set) $Y:=V(\mathcal{I})$ of a ideal-sheaf $\mathcal{I}\subset\mathcal{O}_{X}$ is a subspace of $X$ which is not necessarily a submanifold (it may have singularity as well). The sheaf of functions $\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}}$ on $Y$ is its structure sheaf and the pair $(Y,\mathcal{O}_{Y})$ is an analytic space (see in [Pym13]).
Here, the tangent sheaf $\mathcal{T}_{Y}:=\mathcal{D}er_{\mathbb{C}_{Y}}(\mathcal{O}_{Y})$ is not necessarily be a locally free Lie algebroid over $(Y,\mathcal{O}_{Y})$. Consider the sheaf of logarithmic derivations ((((forms a generalized involutive distribution or singular foliation)))) [Pym13, MS22] as
@@
((((or, sheaf of vector fields on $X$ that are tangent to $Y)$ with the canonical Lie algebroid structure. It is associated with $\mathcal{T}_{Y}$ via the Lie algebroid epimorphism
@@
For example, let $X=\mathbb{C}^{2}$ with global coordinate functions $x,y$ and consider the analytic space as $(Y,\mathcal{O}_{Y})$ associated to the vanishing set of the ideal sheaf $\mathcal{I}$ generated by the function $xy$. Here, the subspace $Y$ is the union of points on the axes, known as a normal crossing divisor, has a singularity at the origin $\bar{0}:=(0,0)$. Thus, the Lie algebroid $\mathcal{T}_{X}(-logY)$ is a locally free $\mathcal{O}_{X}$-module generated by the derivations $x\partial_{x}$ and $y\partial_{y}$, but the Lie algebroid $\mathcal{T}_{Y}$ is not locally free.

### v7

@@ Abstract.
We consider Lie algebroids over an algebraic space (or topological ringed space) as quasicoherent sheaves of Lie-Rinehart algebras. We express hypercohomology for a locally free Lie algebroid (not necessarily of finite rank) as a derived functor, and simplify it via Čech cohomology. Furthermore, we define the Hochschild hypercohomology of a sheaf of generalized bialgebras and study the cases of the universal enveloping algebroid and the jet algebroid of a Lie algebroid. In the sequel, we present a version of Hochschild-Kostant-Rosenberg theorem for a locally free Lie algebroid, as well as its dual version.
@@
AMS Mathematics Subject Classification : $16$S30303030, $16$T10,1010,10 ,$17$B56565656.
@@ 1. Introduction
The notion of Lie algebroids plays a prominent role in differential geometry and mathematical physics as they represent generalized infinitesimal symmetries of spaces, which are related to the corresponding global symmetries of spaces described by Lie groupoids [21]. Lie algebroids over a $C^{\infty}$-manifold are joint generalization of tangent vector bundle over the manifold and Lie algebras [26]. An algebraic analogue of Lie algebroids, known as Lie-Rinehart algebras, is used to study general situations [16, 18, 22, 3]. In the context of complex geometry and algebraic geometry, Lie algebroids over analytic spaces [32] and over algebraic varieties (or schemes) [18, 7, 3] respectively have been studied in sheaf theoretic language, which are joint generalization of the tangent sheaf and sheaves of Lie algebras [18, 33, 32]. We study Lie algebroids over an algebraic space (shortly, $a$-space) or over a topological ringed space $(X,\mathcal{O}_{X})$ as certain quasicoherent sheaves of Lie-Rinehart algebras [37, 29], where $\mathcal{O}_{X}$ is a sub-sheaf of algebras of the sheaf of continuous functions $C_{X}^{0}$ on $X$ [33, 30]. This framework unifies the concept of Lie algebroids across these three types of base spaces as special cases. To study calculus on these (smooth and singular) geometric objects in a consistent manner, we need the notion of Lie algebroids within algebro-geometric settings. It allows one to treat several geometric structures, such as Poisson analytic spaces [32], singular foliations or generalized involutive distributions [10, 32, 20], as well as free Lie algebroids [18]. In particular, the sheaf of logarithmic derivations for some (principal or free) divisor of a complex manifold (or a smooth algebraic variety) [4, 32, 29], and the path algebroid of a smooth manifold [18], are key object of study in this context.
@@
Associated with a Lie algebroid over an $a$-space, there are two canonical sheaves of generalized bialgebras: the universal enveloping algebroid and the jet algebroid [18, 7, 6, 2, 32, 29]. Both play crucial roles in studying the homological algebra of a Lie algebroid. Here, the sheaves of generalized bialgebras serve as a sheaf-theoretic analogue (or global version) of the notion of $R/{\mathbb{K}}$-bialgebras [26] or left $R$-bialgebroids [19], referred to as $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras [29]. The universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ of a Lie algebroid $\mathcal{L}$ generalizes the notion of sheaf of differential operators on a manifold [35, 29]. It is sheafification of the presheaf of universal enveloping algebras of Lie-Rinehart algebras associated with each space of sections of $\mathcal{L}$ [3, 29]. It has a canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structure [29], similar structures are present in [18, 7] by different names. Moreover the dual of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, $\mathscr{J}(\mathcal{O}_{X},\mathcal{L}):=\mathscr{H}om_{\mathcal{O}_{X}}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}),\mathcal{O}_{X})$ is the jet algebroid of $\mathcal{L}$, generalizes the notion of sheaf of jets on a manifold (see[18, 7, 5, 32, 2]). It is sheafification of the presheaf of jet algebras associated with the sheaf of Lie-Rinehart algebras [2, 29]. It also has a canonical $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra structure induced from the structure of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ [29]. Moreover, an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra gives rise to a Lie algebroid over the $a$-space $(X,\mathcal{O}_{X})$, consisting of its sheaf of primitive elements [29].
Cohomology of a Lie algebroid generalizes both Lie algebra cohomology (Chevalley-Eilenberg cohomology) and de Rham cohomology, and is commonly refereed to as Chevalley-Eilenberg-de Rham cohomology [21, 32, 2, 1, 3]. The cohomology of a Lie algebroid $\mathcal{L}$, which is a locally free $\mathcal{O}_{X}$-module of finite rank, over a Noetherian separated scheme [3] or over a complex manifold $(X,\mathcal{O}_{X})$ [32], is described using the derived functor $Ext$. In non-commutative geometry [17, 13, 35], the dual pair consisting of the Hochschild cohomology and homology $(HH^{\bullet}(A),HH_{\bullet}(A))$ of an associative $\mathbb{K}$-algebra $A$ serves as a non-commutative analogue to the classical dual pair $(\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{T}_{X},\wedge^{\bullet}_{\mathcal{O}_{X}}\Omega^{1}_{X})$, which consists of the sheaf of multivector fields and the sheaf of differential forms over a smooth $a$-space $(X,\mathcal{O}_{X})$. This correspondence is given by the standard Hochschild-Kostant-Rosenberg (shortly, HKR) theorem. In [7, 5, 6], the authors establish a version of the HKR theorem, using the dual pair $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$, to explore precalculus up to homotopy for locally free Lie algebroids of finite rank. For that, they introduce an isomorphism of Gerstenhaber algebras, given by a twisted version of the HKR morphism. In this paper, we deduce analogous cohomological results using a different approach to include more general situations, such as relaxing the finite rank condition. To derive such results, we present a global version of cohomological results on Lie-Rinehart algebras [34, 16, 19, 25].
@@
In Section 2, we recall some of the preliminary notions required for the article. In Section 3, we consider Lie algebroid hypercohomology over certain special $a$-spaces. Here, we show that one can define Lie algebroid cohomology, where the underlying $\mathcal{O}_{X}$-module, while locally free, is not necessarily of finite rank, thus encompassing all three geometric setups. The cohomology of a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with coefficient in a $\mathcal{L}$-module $\mathcal{E}$, is given by $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$. Here, we use the result of Lie-Rinehart cohomology as derived functor for a projective Lie-Rinehart algebra [34, 16] as a local description. We apply this result for the sheaf of logarithmic derivations and for $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras. Using the cohomology groups we consider (first) Chern classes [24, 25] in this general set up. Furthermore, we express the Lie algebroid cohomology in terms of Čech cohomology by considering a good open cover [33, 36, 1]. In Section 4, we consider algebraic (analytic) de Rham cohomologies for some free divisors associated with principal ideal sheaves [29] and compute the corresponding logarithmic de Rham cohomologies [4, 27]. We view these cohomologies as Lie algebroid (hyper)cohomology as described in [32, 1, 3].
@@
In the first part of Section 5, we define Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra. As special cases of it, we study Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ associated with a Lie algebroid $\mathcal{L}$. After that, we present a version of Hochschild-Kostant-Rosenbergh (HKR) theorem for locally free Lie algebroids (locally free $\mathcal{O}_{X}$-modules but not necessarily of finite rank) over any of the special $a$-spaces. It provides an isomorphism of graded vector spaces between Hochschild hypercohomology of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and hypercohomology of the sheaf of $\mathcal{L}$-poly vector fields, that is $\mathbb{H}H^{\bullet}(\mathscr{U}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$. This result we get by locally using the algebraic counterpart for projective Lie-Rinehart algebras described in [19]. We derive a result in sheaf cohomology context to simplify the hypercohomology $\mathbb{H}^{\bullet}(X,\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{L})$. Moreover, we present the HKR theorem in some of the special cases. Next we discuss about the dual version of HKR theorem in the generalize setup. Thus we show that for a locally free Lie algebroids $\mathcal{L}$ of finite rank over any of the special $a$-space $(X,\mathcal{O}_{X})$, we get a canonical isomorphism of vector spaces between Hochschild hypercohomology of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ and the Lie algebroid (hyper)cohomology of $\mathcal{L}$, that is $\mathbb{H}H^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))\cong\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{O}_{X})$. This result we get by locally using the algebraic counterpart as described in [19]. Both of these hypercohomologies are obtained by the derived functor $Cotor$. At last, we apply the dual HKR theorem for the tangent sheaf $\mathcal{L}:=\mathcal{T}_{X}$ over non-singular $a$-spaces and obtain some interesting results using the (smooth, analytic or algebraic) de Rham cohomology [8, 36] of $X$.
Acknowledgment. The author acknowledges support from the C. S. I. R. fellowship grant during his PhD at Indian Institute of Technology Kanpur. He wholeheartedly expresses his deepest gratitude to Prof. Ashis Mandal for his immense support and guidance. Additionally, he would like to thank Dr. Satyendra Kumar Mishra for his valuable comments and suggestions, which helped to improve the article.
@@ 2.1. Lie algebroids over algebraic spaces
Here we consider some special locally ringed spaces [38], and then define Lie algebroids over these spaces [29].
@@
Let $X$ be a topological space and $\mathcal{O}_{X}$ a $\mathbb{K}_{X}$-subalgebra of the sheaf of continuous functions $C_{X}^{0}$ on $X$. The pair $(X,\mathcal{O}_{X})$ is said to be an algebraic space, or simply, an $a$-space.
For an $a$-space $(X,\mathcal{O}_{X})$, the sheaf of derivations $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$ of $\mathcal{O}_{X}$, known as the tangent sheaf of $X$.
@@
Consider smooth manifolds, complex manifolds, analytic spaces and algebraic varieties, all with their associated structure sheaf, as special $a$-spaces. The tangent sheaf over such an $a$-space $(X,\mathcal{O}_{X})$ is a quasicoherent module over $\mathcal{O}_{X}$ [18, 32].
@@
A Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ is a quasicoherent sheaf of $(\mathbb{K}_{X},\mathcal{O}_{X})$-Lie-Rinehart algebras. That is, $\mathcal{L}$ is a $\mathbb{K}_{X}$-Lie algebra and a quasicoherent $\mathcal{O}_{X}$-module equipped with a homomorphism $\mathfrak{a}:(\mathcal{L},[\cdot,\cdot])\rightarrow(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c})$ of $\mathcal{O}_{X}$-modules and $\mathbb{K}_{X}$-Lie algebras, called the anchor map. The map $\mathfrak{a}$ satisfies the Leibniz rule: $[D,fD^{\prime}]=f[D,D^{\prime}]+\mathfrak{a}(D)(f)D^{\prime}$ for $f\in\mathcal{O}_{X}$ and $D,D^{\prime}\in\mathcal{L}$.
@@
Lie algebroids over a smooth manifold $X$ is equivalent to locally free Lie algebroids of finite rank over the $a$-space $(X,C_{X}^{\infty})$, where $C_{X}^{\infty}$ is the sheaf of $C^{\infty}$-functions on $X$.
@@
The standard Lie algebroid structure on the tangent sheaf of an $a$-space $(X,\mathcal{O}_{X})$ is described by the pair $(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c},id)=:\mathcal{T}_{X}$. In particular, when $X$ is a real smooth manifold (complex manifold), the sheaf of smooth (holomorphic) vector fields $\mathfrak{X}_{X}$ on $X$ is isomorphic to $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$, forming a locally free Lie algebroid of rank equals to dimension of the manifold $X$.
The sheaf of differential $1$-forms $\Omega^{1}_{X}$ over a Poisson manifold $X$ has a canonical Lie algebroid structure. For smooth and analytic cases the details can be found in [10] and [32] respectively.
@@
A singular foliation $\mathcal{F}$ on a real smooth manifold or a complex manifold $(X,\mathcal{O}_{X})$ is an $\mathcal{O}_{X}$-submodule of the Lie algebroid $\mathcal{T}_{X}$ ((((or $\mathfrak{X}_{X})$, which is $(i)$ stable under the Lie bracket and $(ii)$ locally finitely generated. It provides a Stefan-Sussmann distribution on $X$, forms a Lie algebroid over $(X,\mathcal{O}_{X})$. In particular, regular foliations arise from involutive sub bundles of the tangent bundle ((((providing Frobenius distributions)))). See [10, 32, 20] for details.
@@
Let $X$ be a complex manifold and $\mathcal{O}_{X}$ be the sheaf of holomorphic functions. The vanishing set ((((or the zero locus)))) $Y:=V(\mathcal{I})$ of a ideal-sheaf $\mathcal{I}\subset\mathcal{O}_{X}$ is a subspace of $X$ which is not necessarily a submanifold ((((it may have singularity as well)))). The sheaf of functions $\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}}$ on $Y$ is its structure sheaf and the pair $(Y,\mathcal{O}_{Y})$ is an analytic space (see in [32]).
Here, the tangent sheaf $\mathcal{T}_{Y}:=(\mathcal{D}er_{\mathbb{C}_{Y}}(\mathcal{O}_{Y}),[\cdot,\cdot]_{c},id_{Y})$ is not necessarily a locally free Lie algebroid over $(Y,\mathcal{O}_{Y})$. Consider the sheaf of logarithmic derivations ((((forms a generalized involutive distribution or singular foliation)))) [32, 29] as
@@
((((geometrically, it represents the sheaf of vector fields on $X$ that are tangent to $Y$ for a smooth divisor $Y)$ with the canonical Lie algebroid structure. It is associated with $\mathcal{T}_{Y}$ via the Lie algebroid epimorphism
@@
For example, let $X=\mathbb{C}^{2}$ with global coordinate functions $x,y$ and consider the analytic space as $(Y,\mathcal{O}_{Y})$ associated to the vanishing set of the ideal sheaf $\mathcal{I}$ generated by the function $xy$. Here, the subspace $Y$ is the union of points on the axes, known as a normal crossing divisor, has a singularity at the origin $\bar{0}:=(0,0)$. Thus, the Lie algebroid $\mathcal{T}_{X}(-logY)$ is a locally free $\mathcal{O}_{X}$-module generated by the derivations $x\partial_{x}$ and $y\partial_{y}$, but the Lie algebroid $\mathcal{T}_{Y}$ is not locally free. See [32, 29] for details.